Circular No. 6019

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

PERIODIC COMET SHOEMAKER-LEVY 9 (1993e) T. Rettig, J. Hahn, S. Tegler and G. Sobczak, University of Notre

Dame; and M. Mumma and M. DiSanti, NASA Goddard Space Flight Center, communicate: “Preliminary analysis of images of the cometary fragments (excluding fragments T = 4 and P_{1} = 8a) obtained with the Hubble Space Telescope wide-field camera between January and July shows that the inner 0′′.3 regions have remained relatively constant or have slowly decreased in brightness (by ∼ 0.1–0.4 mag) over the seven-month period. Similar results are noted with diaphragms of diameter 1′′.1 and 2′′.1. The fragments show minimal brightness activity through the May observations. However, for six fragments observed during June and July, we measure a rather dramatic brightness decrease from the May observations. In the 1′′.1 diaphragm, five fragments (W = 1, V = 2, S = 5, R = 6 and H = 14) show a brightness decrease of ∼ 0.8–0.9 mag, and fragment U = 3 shows a brightness decrease of ∼ 0.4 mag. Also in the June and July images, the brightness profiles of fragment 8b = P_{2} show a very much more distributed source of dust within the inner 1′′. These profiles indicate that the central source may have been disrupted and may suggest initial tidal dispersal of the swarm material. We note that between May and June most of the fragment images have become elongated in the direction of Jupiter.”

SUPERNOVA 1994I IN NGC 5194 W. H. G. Lewin, Massachusetts Institute of Technology; H.-U. Zimmer-

mann and W. Pietsch, Max-Planck-Institut für Extraterrestrische Physik; and R. Beck, Max-Planck-Institut für Radio Astronomie, report: “The xray observatory ROSAT observed (with the HRI) SN 1994I in M51 for 9400 s on May 22. SN 1994I is a type Ib/c supernova (IAUC 5964, 5966, 5972, 5978, 5981), and its position is ∼ 18′′ from the x-ray-bright nucleus (IAUC 5961). It was not detected at x-ray wavelengths. Assuming a thermal bremsstrahlung model, an absorbing column density N_{H} = 1.7×10^{20} cm^{−2} and a temperature kT = 5 keV, an upper limit (95.4-percent-confidence level) for the unabsorbed flux from the supernova in the energy range 0.1– 2.4 keV is 2.06 × 10^{−14} erg cm^{−2} s^{−1}. At a distance of 7.7 Mpc, this corresponds to an unabsorbed soft x-ray luminosity of 1.5 × 10^{38} erg/s. The given values are not very sensitive to the assumed temperature (for kT = 1 keV they are about 20 percent lower). Our upper limit on day 53 after outburst is an order of magnitude lower than the detected soft x-ray luminosity of SN 1993J (type IIb) in M81 at the same age.”

1994 July 14 Brian G. Marsden